OpenStudy (anonymous):

How many liters of methane gas (CH4) need to be combusted to produce 12.4 liters of water vapor, if all measurements are taken at the same temperature and pressure? Show all of the work used to solve this problem. CH4 (g) + 2 O2 (g) yields CO2 (g) + 2 H2O (g)

OpenStudy (aaronq):

In stoichiometry, you can use a shortcut when working with gases. Since moles are proportional to volume (when assuming ideal behaviour), simply, use the stoichiometric ratios and the volume of gas. \(\dfrac{V_{CH_4}{1}=\dfrac{V_{H_2O}{2}=\)

OpenStudy (aaronq):

\(\dfrac{V_{CH_4}}{1}=\dfrac{V_{H_2O}}{2}\)

OpenStudy (anonymous):

Thanks! That makes more sense

OpenStudy (aaronq):

no problem !